What rate of change of current in a solenoid having self-inductance 19.4 mH produces a self-induced emf of 100 mV in it?
1
Expert's answer
2020-03-10T11:31:52-0400

electromagnetic law


"\frac{di}{dt}=\frac{\epsilon}{L}"

"\frac{di}{dt}=\frac{100}{19.4}=5 (aps)"